OpenStudy (anonymous):

Solvebequation for x, where 0

jhonyy9 (jhonyy9):

so like a first step 4(cos x)^2 = 3

jhonyy9 (jhonyy9):

do you can continue it ?

OpenStudy (anonymous):

cosx = +&-\[\cos x = \pm \sqrt{3/4}\]

OpenStudy (anonymous):

oh nvm i got it

OpenStudy (anonymous):

thanks

jhonyy9 (jhonyy9):

yes but check it please again and will see it there what can you simplifie there +/- sqrt(3/4) ?

jhonyy9 (jhonyy9):

ok ?